Grade 12th passMechanics

A boat has speed of 5 kilometre per hour in still water crosses a river of width 1 kilometre along the shortest possible path in 15 minutes the velocity of river water in kilometre per hour is first option 1 second option to third option 3 e fourth option under root 41

Sakshi
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago

With no current, the boat would cross the river in 60/5, or 12 minutes. Instead, it takes 15 minutes. In 15 minutes it travels 15/12, or 5/4 kilometers. If this is seen as the hypotenuse of a right triangle, with the long side being 1 kilometer, and the short side being n, we can calculate n by:

(5/4)²-1²=n²

25/16–16/16=9/16

n=3/4 km displacement

3/4 km in 15 minutes=3 km/hr as the speed of the current